# On-Policy Policy Gradient Reinforcement Learning Without On-Policy Sampling

### v2

@@ 5 Proximal Robust On-Policy Sampling for Policy Gradient Algorithms
To address these challenges, we propose a new behavior policy update. Observing that the gradient of the ros loss $\nabla_{\bm{\phi}}{\mathcal{L}}=\nabla_{\bm{\phi}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})\rvert_{{\bm{\phi}}={\bm{\theta}}}$ resembles the policy gradient (Eq. 2) with $A^{\pi_{{\bm{\theta}}}}({\bm{s}},{\bm{a}})=-1,\forall({\bm{s}},{\bm{a}})$, we draw inspiration from ppo [Schulman et al., 2017] to address these challenges. Rather than using the ros objective to update ${\bm{\phi}}$, we use a clipped surrogate objective:
@@
$\begin{split}{\mathcal{L}}_{\textsc{clip}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\textsc{props}}{}{})=\min\Bigg{[}-\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})},-\mathtt{clip}\left(\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})},1-\epsilon_{\textsc{props}}{}{},1+\epsilon_{\textsc{props}}{}{}\right)\Bigg{]}.\end{split}$ (5)
@@
Table 1 in Appendix B summarizes the behavior of ${\mathcal{L}}_{\text{{clip}{}}}$. Intuitively, this objective incentivizes the agent to decrease the probability of observed actions by at most a factor of $1-\epsilon_{\textsc{props}}{}{}$. Let $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})=\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})}$. When $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})<1-\epsilon_{\textsc{props}}{}{}$, this objective is clipped at $-(1-\epsilon_{\textsc{props}}{}{})$. The loss gradient $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$ becomes zero, and the $({\bm{s}},{\bm{a}})$ pair has no effect on the policy update. When $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})>1-\epsilon_{\textsc{props}}{}{}$, clipping does not apply, and the (non-zero) gradient $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$ points in a direction that decreases the probability of $\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$. As in the ppo update, this clipping mechanism avoids destructively large policy updates and permits us to perform many epochs of minibatch updates with the same batch of data.
@@
$\begin{split}{\mathcal{L}}_{\textsc{props}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\textsc{props}}{}{},\lambda)={\mathcal{L}}_{\textsc{clip}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\textsc{props}}{})-\lambda D_{\textsc{kl}}(\pi_{{\bm{\theta}}}(\cdot|{\bm{s}})||\pi_{\bm{\phi}}(\cdot|{\bm{s}}))\end{split}$ (6)
where $\lambda$ is a regularization coefficient quantifying a trade-off between maximizing ${\mathcal{L}}_{\text{{props}{}}}$ and minimizing $D_{\textsc{kl}}$. We provide full pseudocode for the props update in Algorithm 2. Like ros, we set the behavior policy parameters ${\bm{\phi}}$ equal to the target policy parameters at the start of each behavior update, and then make a local adjustment to $\phi$ to increase the probabilities of under-sampled actions. We stop the props update when $D_{\textsc{kl}}{}(\pi_{\bm{\theta}}||\pi_{\bm{\phi}})$ reaches a chosen threshold $\delta_{\textsc{props}}{}$. This technique further safeguards against large policy updates and is used in widely adopted implementations of ppo [Raffin et al., 2021, Liang et al., 2018]. The props update allows us to efficiently learn a behavior policy that keeps the distribution of data in the buffer close to the expected distribution of the target policy.
@@ 6 Experiments
The central goal of our work is to understand whether on-policy policy gradient algorithms are more data efficient learners when they use on-policy data acquired without on-policy sampling. Towards this goal, we design experiments to answer the two questions:
@@ 1. Q1:
    Does props achieve lower sampling error than on-policy sampling during policy gradient learning (i.e., does props produce data that is more on-policy than on-policy sampling produces)?
@@ 2. Q2:
    Does props increase the fraction of training runs that converge to high-return policies and improve the data efficiency of on-policy policy gradient algorithms?
Our empirical analysis focuses on continuous-state continuous-action MuJoCo benchmark tasks and a tabular 5x5 GridWorld task (Figure 2(a)). We additionally consider three continuous-state discrete-action tasks: CartPole-v1, LunarLander-v2, and Discrete2D100-v0 – a 2D navigation task with 100 discrete actions. Due to space constraints, we include these tasks in Appendix D.4.
@@
(c) Cosine similarity
Figure 2: (a) A GridWorld task in which the agent receives reward $+1$ upon reaching the bottom right corner (the optimal goal), a reward of $+0.5$ upon reaching the top left corner (the suboptimal goal), and a reward of $-0.01$. The agent always starts in the center of the grid. Under an initially uniform policy, the agent visits both goals with equal probability, and thus the true policy gradient increases the probability of reaching the optimal goal. (b, c) Samping error and gradient cosine similarity over 10 seeds. props reduces sampling error and achieves more accurate gradients faster than on-policy sampling. Shaded regions denote 95% bootstrap confidence intervals.
@@ 6.1 Correcting Sampling Error for a Fixed Target Policy
We first study how quickly props decreases sampling error when the target policy is fixed. This setting is similar to the policy evaluation setting considered by  Zhong et al. [2022]. As such, we provide two baselines for comparison: on-policy sampling and ros.
Sampling error metrics. In GridWorld, we compute sampling error as the total variation (TV) distance between the empirical state-action visitation $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$ distribution – denoting the proportion of times $({\bm{s}},{\bm{a}})$ appears in buffer ${\mathcal{D}}$ – and the true state-action visitation distribution under the agent’s policy: $\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}|d_{\mathcal{D}}({\bm{s}},{\bm{a}})-d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})|$. In continuous MuJoCo tasks where it is difficult to compute $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$, we follow Zhong et al. [2022] and measure sampling error using the kl-divergence $D_{\textsc{kl}{}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})$ between the empirical policy $\pi_{\mathcal{D}}$ and the target policy $\pi_{\bm{\theta}}$. We estimate $\pi_{\mathcal{D}}$ as the maximum likelihood estimate under data in the buffer via stochastic gradient ascent. Further details on how we compute $\pi_{\mathcal{D}}$ are in Appendix C.
Since it is straightforward to compute the true policy gradient in the GridWorld task, we additionally investigate how sampling error reduction affects gradient estimation by measuring the cosine similarity between the empirical policy gradient $\nabla_{\bm{\theta}}\widehat{J}({\bm{\theta}})$ and the true policy gradient. As the empirical gradient aligns more closely with the true gradient, the cosine similarity approaches 1.
(a) Fixed target policy.
(b) During rl training.
Figure 3: props reduces sampling error faster than on-policy sampling and ros. In (a), the ros and on-policy sampling curves overlap. Solid curves denote means over 5 seeds. Shaded regions denote 95% confidence intervals.
Experimental setup. In all tasks, we use a buffer with capacity of $\lfloor T/2\rfloor$ samples, where $T$ is the total number of samples collected by the agent. Thus, we expect sampling error to decrease over the first $\lfloor T/2\rfloor$ samples and then remain roughly constant afterwards once the buffer is full. We use randomly initialized target policies. Further experimental details such as hyperparameter tuning are described in Appendix D.2.
Results. As shown in Figure 2(b) and 2(c), in GridWorld, props decreases sampling error faster than on-policy sampling, resulting in more accurate policy gradient estimates. props and ros perform similarly, though this behavior is expected; in a tabular setting with a fixed target policy (i.e. there is no off-policy data in the buffer), we do not encounter Challenge 1 and 2 described in the previous section. In continuous MuJoCo tasks where Challenge 2 arises, props decreases sampling error faster than on-policy sampling and ros (Figure 3(a)). In fact, ros shows little to no improvement over on-policy sampling in every MuJoCo task. This limitation of ros matches the findings of Zhong et al. [2022] where ros struggled to reduce sampling error even in low-dimensional continuous-action tasks. Moreover, props decreases sampling error without clipping and regularization, emphasizing how adaptive off-policy sampling alone decreases sampling error. Due to space constraints, we include results for the remaining environments in Appendix D.2. We additionally include experiments using a fixed, randomly initialized target policy as well as ablation studies isolating the effects of props’s objective clipping and regularization in Appendix D.2. Results with a random target policy are qualitatively similar to those in Figure 3(a), and we observe that clipping and regularization both individually help reduce sampling error. In Appendix D.1, we show empirically that the empirical state-action visitation distributions produced by props (and ros) are unbiased and lower variance than the distribution produced by on-policy sampling.
@@ 6.2 Correcting Sampling Error During RL Training
We are ultimately interested in understanding how replacing on-policy sampling with props affects the data efficiency of on-policy learning, where the target policy is continually updated. In the following experiments, we train rl agents with props and on-policy sampling to evaluate (1) the data efficiency of training, (2) the distribution of returns achieved at the end of training, and (3) the sampling error throughout training. We use the same sampling error metrics described in the previous section and measure data efficiency as the return achieved within a fixed training budget. Since ros [Zhong et al., 2022] is computationally expensive and fails to reduce sampling error in MuJoCo tasks even with a fixed policy, we omit it from MuJoCo experiments.
@@
Experimental setup. We use ppo [Schulman et al., 2017] to update the target policy. We consider two baseline methods for providing data to compute ppo updates: (1) vanilla ppo with on-policy sampling, and (2) ppo with on-policy sampling and a buffer of size $b$ (ppo-buffer). ppo-buffer is a naive method for improving data efficiency of on-policy algorithms by reusing off-policy data collected by old target policies as if it were on-policy data. Although ppo-buffer computes biased gradients, it has been successfully applied in difficult learning tasks [Berner et al., 2019]. Since props and ppo-buffer have access to the same amount of data for each policy update, any performance difference between these two methods can be attributed to differences in how they sample actions during data collection.
@@
In MuJoCo experiments, we set $b=2$ such that agents retain a batch of data for one extra iteration before discarding it. In GridWorld, we use $b=1$ and discard all historic data. Since props and ppo-buffer compute target policy updates with $b$ times as much learning data as ppo, we integrate this extra data by increasing the minibatch size for target and behavior policy updates by a factor of $b$. Further experimental details including hyperparameter tuning are described in Appendix D.6. For MuJoCo tasks, we plot the interquartile mean (IQM) return throughout training as well as the distribution of returns achieved at the end of training (i.e., the performance profile) [Agarwal et al., 2021]. For GridWorld, we plot the agent’s success rate, the fraction of times it finds the optimal goal.
(a) Training curves.
(b) Performance profiles.
Figure 4: (a) IQM returns over 50 seeds. Shaded regions denote 95% bootstrap confidence intervals. (b) Performance profiles over 50 seeds. Higher values correspond to more reliable convergence to high-return policies. Shaded regions denote 95% bootstrap confidence intervals.
Results. As shown in Figure 5(a), on-policy sampling has approximately a 77% success rate on GridWorld, whereas props and ros achieve 100% success rate. In Figure 4(a), props achieves higher return than both ppo and ppo-buffer throughout training in all MuJoCo tasks except Ant-v4 where props dips slightly below ppo’s return near the end of training. Moreover, in Figure 4(b), the performance profile of props almost always lies above the performance profiles of ppo and ppo-buffer, indicating that any given run of props is more like to obtain a higher return than ppo-buffer. Thus, we affirmatively answer Q2 posed at the start of this section: props increases the fraction of training runs with high return and increases data efficiency.
@@
In Appendix D.4, we provide additional experiments demonstrating that props improves data efficiency in discrete-action tasks. We additionally ablate the buffer size $b$ in Appendix D.3. We find that data efficiency may decrease with a larger buffer size. Intuitively, the more historic data kept around, the more data that must be collected to impact the aggregate data distribution.
@@
As shown in Figure 5(b), props achieves lower sampling error than on-policy sampling with a buffer in Humanoid-v4. Due to space constraints, we provide sampling error curves for the remaining MuJoCo environments in Appendix D.3. In GridWorld, props and ros reduce sampling error in the first 300 steps and closely matches on-policy sampling afterwards. We use a batch size of 80 in these experiments, and as the target policy becomes more deterministic, larger batch sizes are needed to observe differences between props and on-policy sampling. [^2] When the target policy is deterministic, we always have zero sampling error, and props will exactly match on-policy sampling.
We additionally ablate the effects of the clipping coefficient $\epsilon_{\textsc{props}}{}{}$ and regularization coefficient $\lambda$ in Appendix D.3. Without clipping or without regularization, props often achieves greater sampling error than on-policy sampling, indicating that both help to keep sampling error low. Moreover, data efficiency generally decreases when we remove clipping or regularization, showing both are essential to props. Thus, we affirmatively answer Q1 posed at the start of this section: props achieves lower sampling error than on-policy sampling when the target policy is fixed and during rl training.

### v3

@@ 5.2 Proximal Robust On-Policy Sampling
To address the challenges ROS faces in continuous control, we propose a new behavior policy update. To address Challenge 1, first observe that the gradient of the ROS loss $\nabla_{\bm{\phi}}{\mathcal{L}}=\nabla_{\bm{\phi}}\log\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})\rvert_{{\bm{\phi}}={\bm{\theta}}}$ is equivalent to the policy gradient (Eq. 2) with $A^{\pi_{{\bm{\theta}}}}({\bm{s}},{\bm{a}})=-1,\forall({\bm{s}},{\bm{a}})$. Since the clipped surrogate objective of PPO (Eq. 3) prevents destructively large updates in on-policy policy gradient learning, we can use a similar clipped objective in place of the ROS objective to prevent destructive behavior policy updates:
@@
$\begin{split}{\mathcal{L}}_{\textsc{clip}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\text{PROPS}}{})=\min\Bigg[-\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})},-\mathtt{clip}\left(\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})},1-\epsilon_{\text{PROPS}}{},1+\epsilon_{\text{PROPS}}{}\right)\Bigg].\end{split}$ (8)
@@
Table 1 in Appendix C summarizes the behavior of ${\mathcal{L}}_{\text{{clip}{}}}$. Intuitively, this objective is equivalent to the PPO objective (Eq. 3) with $A({\bm{s}},{\bm{a}})=-1,\forall({\bm{s}},{\bm{a}})$ and incentivizes the agent to decrease the probability of observed actions by at most a factor of $1-\epsilon_{\text{PROPS}}{}$. Let $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})=\frac{\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})}{\pi_{{\bm{\theta}}}({\bm{a}}|{\bm{s}})}$. When $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})<1-\epsilon_{\text{PROPS}}{}$, this objective is clipped at $(1-\epsilon_{\text{PROPS}}{})$. The loss gradient $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$ becomes zero, and the $({\bm{s}},{\bm{a}})$ pair has no effect on the policy update. When $g({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})>1-\epsilon_{\text{PROPS}}{}$, clipping does not apply, and the gradient $\nabla_{\bm{\phi}}{\mathcal{L}}_{\text{{clip}{}}}$ points in a direction that decreases the probability of $\pi_{\bm{\phi}}({\bm{a}}|{\bm{s}})$. As in the PPO update, this clipping mechanism avoids destructively large policy updates and permits us to perform multiple epochs of minibatch updates with the same batch of data.
@@
$\begin{split}{\mathcal{L}}_{\text{PROPS}}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}},\epsilon_{\text{PROPS}}{},\lambda)={\mathcal{L}}_{\textsc{clip}}{}({\bm{s}},{\bm{a}},{\bm{\phi}},{\bm{\theta}})-\lambda D_{\text{KL}}(\pi_{{\bm{\theta}}}(\cdot|{\bm{s}})||\pi_{\bm{\phi}}(\cdot|{\bm{s}}))\end{split}$ (9)
where $\lambda$ is a regularization coefficient quantifying a trade-off between maximizing ${\mathcal{L}}_{\text{PROPS}}$ and minimizing $D_{\textsc{kl}}$. Algorithm 2 provides pseudocode for the PROPS update. Like ROS, we set the behavior policy parameters $\bm{\phi}$} equal to the target policy parameters at the start of each behavior update, and then make a local adjustment to $\phi$ to increase the probabilities of under-sampled actions. We stop the PROPS update early when $D_{\text{KL}}(\pi_{\bm{\theta}}||\pi_{\bm{\phi}})$ reaches a chosen threshold $\delta_{\text{PROPS}}$. This technique further safeguards against large policy updates and is used in widely adopted implementations of PPO (Raffin et al., 2021; Liang et al., 2018). [^4] The KL regularizer Eq. 9 and the KL cutoff rule $D_{KL}(\pi_{\theta}||\pi_{\phi})>\delta_{\text{PROPS}}$ serve different purposes. KL regularization limits how far the behavior policy can move from the target policy after each gradient step, while the cutoff rule prevents the behavior policy from drifting too far across multiple epochs of minibatch updates, even if each step individually is small In Appendix B, we provide theoretical intuition for the relationship between different PROPS hyperparameters. PROPS enables us to efficiently learn a behavior policy that keeps the distribution of data in the buffer close to the expected distribution of the target policy.
@@ 6 Experiments
The central goal of our work is to understand if reducing sampling error with adaptive, off-policy sampling results in more data-efficient on-policy policy gradient learning. Towards this goal, we design experiments on continuous-state and continuous-action MuJoCo benchmark tasks (Brockman et al., 2016) and a tabular 5x5 GridWorld task (Fig. 2(a)) to answer the following questions:
@@ 1. Q1:
    Does PROPS achieve lower sampling error than on-policy sampling during RL training?
@@ 2. Q2:
    Does PROPS improve the data efficiency of on-policy policy gradient algorithms? That is, does PROPS reduce the number of environment interactions required to achieve returns comparable to on-policy sampling?
6.1 Sampling Error Metrics
In GridWorld, we compute sampling error as the total variation (TV) distance between the empirical state-action visitation $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$ distribution—denoting the proportion of times $({\bm{s}},{\bm{a}})$ appears in buffer $\mathcal{D}$}—and the true state-action visitation distribution under the agent’s policy: $\sum_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}|d_{\pi_{\bm{\theta}}}({\bm{s}},{\bm{a}})-d_{\mathcal{D}}({\bm{s}},{\bm{a}})|$. Since it is straightforward to compute the true policy gradient in the GridWorld task, we additionally investigate how sampling error reduction affects gradient estimation by measuring the cosine similarity between the empirical policy gradient $\nabla_{\bm{\theta}}\widehat{J}({\bm{\theta}})$ and the true policy gradient. As the empirical gradient aligns more closely with the true gradient, the cosine similarity approaches 1. In continuous MuJoCo tasks where it is difficult to compute $d_{\mathcal{D}}({\bm{s}},{\bm{a}})$, we compute sampling error as the KL-divergence $D_{\text{KL}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})$ between the empirical policy $\pi_{\mathcal{D}}$ and the target policy $\pi_{\bm{\theta}}$, which is the primary metric Zhong et al. (2022) used to show ROS reduces sampling error:
$D_{\text{KL}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})=\mathbb{E}_{{\bm{s}}\sim{\mathcal{D}},{\bm{a}}\sim\pi_{\mathcal{D}}(\cdot|{\bm{s}})}\left[\log\left(\frac{\pi_{\mathcal{D}}({\bm{a}}|{\bm{s}})}{\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})}\right)\right].$ (10)
We estimate $\pi_{\mathcal{D}}$ as the maximum likelihood estimate under data in the buffer via stochastic gradient ascent. More concretely, we let $\bm{\theta}$$^{\prime}$ be the parameters of a neural network with the same architecture as $\pi_{\bm{\theta}}$ and then compute:
${\bm{\theta}}_{\textsc{MLE}}=\operatorname*{arg\,max}_{{\bm{\theta}}^{\prime}}\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\log\pi_{{\bm{\theta}}^{\prime}}({\bm{a}}|{\bm{s}})$ (11)
$D_{\text{KL}}(\pi_{\mathcal{D}}||\pi_{\bm{\theta}})\approx\sum\nolimits_{({\bm{s}},{\bm{a}})\in{\mathcal{D}}}\left(\log\pi_{{\bm{\theta}}_{\textsc{MLE}}}({\bm{a}}|{\bm{s}})-\log\pi_{\bm{\theta}}({\bm{a}}|{\bm{s}})\right).$ (12)
@@ 6.1 Sampling Error Metrics
(c) Gradient accuracy
Figure 2: (a) A GridWorld task in which the agent receives reward +1+1 upon reaching the bottom right corner (the optimal goal), a reward of +0.5+0.5 upon reaching the top left corner (the suboptimal goal), and a reward of 0.01-0.01. The agent always starts in the center of the grid. Under an initially uniform policy, the agent visits both goals with equal probability, and thus the true policy gradient increases the probability of reaching the optimal goal. However, sampling error can yield an empirical gradient that increases the probability of reaching the suboptimal goal and cause the agent to converge suboptimally. To converge optimally, the agent must have low sampling error. (b, c) PROPS reduces sampling error and achieves more accurate gradients faster than on-policy sampling. Solid curves denote means over 50 seeds. Shaded regions denote 95% bootstrap confidence belts.
@@ 6.2 Correcting Sampling Error for a Fixed Target Policy
We first study how quickly PROPS decreases sampling error when the target policy is fixed. This setting is similar to the policy evaluation setting considered by  Zhong et al. (2022). As such, we provide two core baselines for comparison: on-policy sampling and ROS. In GridWorld, we additionally include an oracle sampling method that always selects the most under-sampled state–action pair $({\bm{s}}^{*},{\bm{a}}^{*})=\arg\max_{({\bm{s}},{\bm{a}})}d_{\pi_{\theta}}({\bm{s}},{\bm{a}})-d_{\mathcal{D}}({\bm{s}},{\bm{a}})$. This sampler is not realizable in standard RL settings because it assumes the ability to reset the agent to an arbitrary state, but it provides a useful reference point for the minimum sampling error achievable for a given number of samples. We provide details on hyperparameter tuning in Appendix D.2.
Results. We collect 8192 samples with a randomly initialized policy and compute sampling error throughout data collection. As shown in Fig. 2(b) and 2(c), in GridWorld, PROPS decreases sampling error faster than on-policy sampling, resulting in more accurate policy gradient estimates. PROPS and ROS perform similarly, though this behavior is expected: in a tabular setting with a fixed target policy, there is no off-policy data in the buffer, so we do not encounter Challenge 1 and 2 described in the previous section. In Appendix D.1, we empirically demonstrate that PROPS is unbiased and has lower variance than on-policy sampling. The oracle sampler attains lower sampling error and higher gradient accuracy than PROPS, as expected, since it assumes access to state resets. [^5] Sampling error does not decrease monotonically under the oracle sampler. With $T$ samples, the expected number of visits to each $({\bm{s}},{\bm{a}})$ under on-policy sampling is $T\cdot d_{\pi_{\theta}}({\bm{s}},{\bm{a}})$, but these quantities are generally not integers, so the oracle sampler must therefore choose the nearest integer allocation, which forces some pairs to be over- or under-sampled. Depending on $T$, this rounding may align more or less closely with the target distribution, leading to fluctuations in the minimum achievable TV distance. For a concrete illustration, suppose there are $n$ state–action pairs and $d_{\pi_{\theta}}$ is uniform. When $T$ is a multiple of $n$, the oracle can allocate counts evenly across all pairs; for other values (e.g., $n+1,\dots,2n-1$), at least one pair must receive an extra or missing count, necessarily increasing sampling error.
In continuous MuJoCo tasks where Challenge 2 arises, PROPS decreases sampling error faster than on-policy sampling and ROS (Fig. 3). In fact, ROS shows no improvement over on-policy sampling in all MuJoCo tasks. This limitation of ROS is unsurprising, as Zhong et al. (2022) showed that ROS struggled to reduce sampling error even in low-dimensional continuous-action tasks. Since batch sizes in the range of 2048–8192 are standard for RL in these tasks, ROS alone will not be able to improve data efficiency during RL. In Appendix D.2, we include similar experiments with expert target policies as well as ablations on PROPS’s objective clipping and regularization. Results with expert target policies are qualitatively similar to Fig. 3, and we observe that clipping and regularization both individually help reduce sampling error.
@@ 6.3 Correcting Sampling Error During RL Training
We are ultimately interested in understanding how replacing on-policy sampling with PROPS affects the data efficiency of on-policy learning. In the following experiments, we train RL agents with PROPS and on-policy sampling to evaluate (1) the data efficiency of training, (2) the distribution of returns achieved at the end of training, and (3) the sampling error throughout training. We use the same sampling error metrics described in the previous section and measure data efficiency as the return achieved within a fixed training budget. Since ROS is computationally expensive and fails to reduce sampling error in MuJoCo tasks even with a fixed policy, we omit it from MuJoCo experiments.
@@
Experimental setup. We use PPO (Schulman et al., 2017) to update the target policy and consider two baseline methods for providing data to compute PPO updates: (1) vanilla PPO with on-policy sampling, and (2) PPO with on-policy sampling using a buffer of size $b$ (PPO-Buffer). PPO-Buffer is a naive method for improving data efficiency of on-policy algorithms by reusing off-policy data collected by old target policies as if it were on-policy data. Although PPO-Buffer computes biased gradients, it has been successfully applied in difficult learning tasks (Berner et al., 2019). Since PROPS and PPO-Buffer have access to the same amount of data for each policy update, any performance difference between these two methods arises from differences in how they sample actions during data collection.
@@
In MuJoCo experiments, we set $b=2$ such that agents retain each batch of data for one extra iteration before discarding it. In GridWorld, we use $b=1$ and discard all historic data. Since PROPS and PPO-Buffer compute target policy updates with $b$ times as much learning data as PPO, we integrate this extra data by increasing the minibatch size for target and behavior policy updates by a factor of $b$. Further experimental details including hyperparameter tuning are described in Appendix F. For MuJoCo tasks, we plot the mean return throughout training as well as the distribution of returns achieved at the end of training (i.e., the performance profile) (Agarwal et al., 2021). For GridWorld, we plot the fraction of times agents find the optimal goal (i.e., the success rate).
(a) Mean normalized return (aggregated).
(b) Performance profiles (aggregated).
Figure 4: Mean normalized return and performance profiles aggregated over all six MuJoCo tasks. We compute normalized returns as $\frac{R_{\text{max}}-R_{t}}{R_{\text{max}}}$, where $R_{\text{max}}$ is the maximum return achieved by any algorithm in a particular task, and $R_{t}$ is the return at timestep $t$. Solid curves denote the mean over 50 seeds per task (300 seeds total). Shaded regions denote 95% bootstrap confidence belts.
@@
Results. Fig. 5(a) shows that on-policy sampling has approximately a 77% success rate on GridWorld, whereas PROPS and ROS achieve a 100% success rate. Fig. LABEL:fig:data_eff and Table LABEL:tab:returns show that PROPS can achieve the same returns as PPO and PPO-Buffer with fewer environment interactions. For instance, in Humanoid-v4 and Hopper-v4, PROPS closely matches the final return of PPO with 33% fewer interactions. To provide a more compact and readable summary of performance throughout training, we also report results aggregated across all six MuJoCo tasks in Fig. 4(a). PROPS matches the normalized return of PPO and PPO-Buffer using only 60% of the training budget. Moreover, the performance profiles of PROPS in Fig. LABEL:fig:performance_profile and Fig. 4(b) are almost always above the performance profiles of PPO and PPO-Buffer, indicating that any given run of PROPS is more likely to obtain a higher return than PPO-Buffer. Thus, we affirmatively answer Q2 posed at the start of this section: PROPS increases the fraction of training runs with high return and increases data efficiency.
@@
As shown in Fig. 6, PROPS achieves lower sampling error than on-policy sampling throughout RL training in 5 out of 6 tasks. In HalfCheetah-v4, PROPS decreases sampling error only in the first 400k timesteps, but nevertheless improves data efficiency. This result likely reflects our hyperparameter tuning procedure in which we selected hyperparameters yielding the largest return (Appendix F). Although lower sampling error intuitively correlates with increased data efficiency, it is nevertheless possible to achieve high return without reducing sampling error. Alternatively, the early reduction in sampling error itself may be sufficient to explain the observed efficiency gains. In GridWorld, PROPS and ROS reduce sampling error in the first 300 steps and closely match on-policy sampling afterwards. We use a batch size of 80 in these experiments, and as the target policy becomes more deterministic, larger batch sizes are needed to observe differences between PROPS and on-policy sampling. [^6] As a policy becomes deterministic, sampling error approaches zero, so there is less sampling error for PROPS to correct.
We ablate the effects of the clipping coefficient $\epsilon_{\text{PROPS}}{}$, regularization coefficient $\lambda$, and buffer size $b$ in Appendix D.3. Without clipping or without regularization, PROPS often achieves greater sampling error than on-policy sampling, indicating that both help to keep sampling error low. Moreover, data efficiency generally decreases when we remove clipping or regularization, showing both are essential to PROPS. We find that data efficiency may decrease with a larger buffer size. Intuitively, the more historic data kept around, the more data that must be collected to impact the aggregate data distribution. Last, we show that PROPS is robust to hyperparameter choices using a sensitivity analysis in Appendix G. Thus, we affirmatively answer Q1 posed at the start of this section: PROPS achieves lower sampling error than on-policy sampling when the target policy is fixed and during RL training.